OpenStudy (anonymous):

lim x-> infinite rad(x^2 +5)/x-2 Evaluate the limit

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\left(\begin{matrix}\sqrt{x ^{2}+5} \\ x-2\end{matrix}\right)\]

OpenStudy (zarkon):

use \frac{\sqrt{x^2+5}}{x-2} \[\frac{\sqrt{x^2+5}}{x-2}\]

OpenStudy (zarkon):

\[\frac{\sqrt{x^2+5}}{x-2}\] \[=\frac{\sqrt{x^2(1+5/x^2)}}{x(1-2/x)}\] \[=\frac{x\sqrt{1+5/x^2}}{x(1-2/x)}\] \[=\frac{\sqrt{1+5/x^2}}{1-2/x}\] now take limit

OpenStudy (anonymous):

How did you get 5/x^2?

OpenStudy (anonymous):

oh you factored

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

Okay! I got it! Thanks :)